OpenStudy (anonymous):

Which of the following is the slope between the two points (0, -4) and (-2, 5)? -nine halves two ninths -two ninths nine halves

OpenStudy (anonymous):

If you follow my link, click on the very first link in the search and read the article, you will never ask a similar question.

OpenStudy (texaschic101):

formula : slope (m) = (y2 - y1) / (x2 - x1) set 1 (0,-4)......set 2 (-2,5) m = (5 - (-4) / (-2 - 0) m = (5 + 4) / -2 m = 9/-2 or - 9/2 (- nine halves) Any questions ?

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

help with another?

OpenStudy (texaschic101):

possibly...math is not my best subject, but let me see what you have

OpenStudy (anonymous):

You should not provide answers to people who are not willing to put in the effort themselves. He will not learn by copying down your working out without understanding anything.

OpenStudy (texaschic101):

your right...my bad

OpenStudy (texaschic101):

I sometimes like to give people a little break....homework sucks

OpenStudy (texaschic101):

you got another problem @JOshua_Lewis

OpenStudy (anonymous):

well @Bad2zBone im on a 500 question test and only havent done not even a portion. i have been up all freaking night doing the test. So excuse me if i would like some help. and i have to pass this if i wanna go on my mission trip to africa

OpenStudy (anonymous):

@terenzreignz if you dont mind then yes :)

OpenStudy (anonymous):

Is you are that bothered about just 'passing the test' without actually having the knowledge to pass the test, pay some money, i could get you at least 95% on this test.

OpenStudy (texaschic101):

and you are talking to me about giving answers....lol

OpenStudy (anonymous):

ikr! lol

OpenStudy (anonymous):

pay some money my ***

OpenStudy (anonymous):

Consider the line that passes through the points (-5, 9) and (-2, 0). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 9 = -1 third(x + 5) y - 9 = -3(x + 5) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (texaschic101):

m = (y2 - y1) / (x2 - x1) set 1 (-5,9).....set 2 (-2,0) m = (0 - 9) / (-2 - (-5) m = (-9)/3 m = - 3 y = mx + b using points (-5,9) 9 = -3(-5) + b 9 = 15 + b 9 - 15 = b - 6 = b y = -3x -6 I think the answer is the last one....neither one represent the line that passes through the two given points

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4

OpenStudy (texaschic101):

y = - 1/2x + b now sub points (4,-6) in -6 = -1/2 (4) + b -6 = -2 + b -6 + 2 = b -4 = b y = - 1/2x - 4 <-- answer

OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y - 4 = 3(x - 2) y - 2 = 3(x - 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

OpenStudy (texaschic101):

y - y1 = m(x - x1).....slope(m) = 3....points (2,4) y - 4 = 3(x - 2)

OpenStudy (anonymous):

Choose the equation of the vertical line that passes through the point (-2, -1). x = -2 y = -1 y = -2 x = -1

OpenStudy (anonymous):

im sorry im not to good at algebra :/

OpenStudy (texaschic101):

x = -2

OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 4y = 8 4y = -2x + 8 y = -one halfx + 2 x = -2y + 4

OpenStudy (texaschic101):

the slope intercept form is y = mx + b. You can probably do this one. Keep in mind, y has to be by itself no numbers attached to it. So what do you think the answer is ?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

c?

OpenStudy (texaschic101):

you got it...very good

OpenStudy (texaschic101):

I am sorry, but I got to go now.....good luck with the rest. I hope you get to go to Africa. Watch out for the lions

OpenStudy (anonymous):

lolol thnaks :)